OpenStudy (anonymous):

PLZ really need help! Simplify the expression. √27x^5 + x√3 + 9√3x If the simplified expression is written in standard form, what is the leading coefficient?

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

Let g and f be two functions defined by g(x)=5-4x^2 f(x)=9-5x for any real number . Find g/f(-5) and then find all values that are NOT in the domain of g/f . If there is more than one value, separate them with commas.

OpenStudy (anonymous):

plz help when you are done ofcourse

OpenStudy (anonymous):

i can't do it lol my brain is friend could you explain a bit more